# Orientation reconstruction of transformation titanium alloys via polarized light microscopy: methodology and assessment

### v1

@@ 1 Introduction
Research into fundamental crystal-scale behavior has led to a decrease in uncertainty with respect to macroscopic mechanical behavior of polycrystalline materials, which has in turn facilitated the design of materials and engineering components which are able to meet increasingly stringent demands in modern machinery. Significant experimental developments were made in the latter half of the 20th century regarding a deeper understanding of how the morphology of grains and the collective distribution of crystallographic orientations (or: crystallographic texture) influence the behavior of materials, primarily in terms of the yield strength and ductility. In parallel, developments in microstructurally-sensitive deformation modeling techniques have similarly enjoyed significant progress. Various techniques have been developed in this regard, from mean field methods in the 1930s (Sachs, 1928; Taylor, 1938), to self-consistent modeling primarily developed in the 1980s and 1990s (Lebensohn and Tomé, 1993). First developed in the late 1980s (Asaro and Needleman, 1985; Nemat-Nasser and Obata, 1986) as a method for modeling texture evolution, crystal plasticity finite element modeling (CPFEM) has evolved to a full-field method capable of capturing the intragrain deformation response of high-fidelity representations of microstructures (considering both crystallographic texture and explicit representations of complex grain morphology) (Marin and Dawson, 1998a, b; Roters et al., 2010), and has been proven adept at the prediction of both intragrain and macroscopic behavior, and has offered unparalleled insight into the relationship of elasticity and plasticity at those scales (and thus our best window into structure-property relationships).
@@
In all of the above modeling methods, the models require some degree of experimental material characterization or micromechanical testing to aid in initiation, verification, or validation of the model. For CPFEM specifically, proper instantiation of a virtual polycrystalline sample that faithfully represents a physical sample requires (at least) information about the geometric morpohology of the grains, as well as the grains orientations (i.e., crystallographic texture) (Roters et al., 2010). Most often, two-dimensional characterization methods such as electron backscatter diffraction (EBSD) are used to measure both granular morphology and orientations on the surfaces of samples, and various methods have been developed to infer the three-dimensional morphology (either statistically (Takahashi and Suito, 2003; Javaheri and Sundararaghavan, 2020), based on oblique surface measurements (Turner and Kalidindi, 2016), or destructively in the case of serial sectioning (Echlin et al., 2020)). More recently, high energy diffraction microscopy (HEDM) has offered the ability to non-destructively map orientations in three dimensions (Suter et al., 2006; Bernier et al., 2011; Pagan et al., 2018; Nygren et al., 2020; Pagan et al., 2021), which can be used to directly construct digital twins of actual physical samples (Quey and Renversade, 2018), rather than statistically-representative virtual samples. Further, in the case of HEDM, micromechanical testing may be performed on the samples after initial characterization, allowing for unprecedented experimental deduction of structure-property relationships (Hurley and Pagan, 2019; Chatterjee et al., 2019).
@@
Unfortunately, while EBSD and HEDM have become well-accepted as characterization techniques and are indeed powerful tools, they suffer from practical limitations. Both are relatively expensive, time consuming, and require a relatively high degree of training necessary to produce skilled researchers capable of performing proper, meaningful measurements. Consequently, it is beneficial to develop methods which significantly reduce cost, lead-time, and/or the knowledge/skill barrier necessary to produce adequate characterization results. Recently, significant research attention has been given to the development of ultrasonic (Lan et al., 2018; Dryburgh et al., 2020; Liu et al., 2021; He et al., 2022) and optical techniques in an effort to address these problems. Of interest to this study, we highlight the advent of polarized light microscopy (PLM) as a characterization method for materials comprised of crystals exhibiting hexagonal symmetry (Jin and De Graef, 2020; Hoover et al., 2019). PLM offers the ability to measure spatial fields of (partial) crystal orientations both relatively rapidly and using relatively affordable equipment (compared to EBSD), as well as with fewer material preparation steps. However, PLM is limited in that it only measures an incomplete description of the orientation of HCP crystals, which limits its applicability as a characterization method for micromechanical simulations owing to the potential for high errors in predictions (van Wees et al., 2023).
@@
In this study, we outline a method to reconstruct full orientations of the crystals in titanium materials to one of four orientations—i.e., a quasi-deterministic approach. As prime inspiration for this study, we consult the work of Glavicic et. al. (Glavicic et al., 2003), which established a method exploiting the physics of the allotropic phase transformation that many titanium alloys undergo during cooling in an effort to deduce the orientation field of the BCC ( phase) microstructure that likely existed at high temperature based on the measurement of the room temperature phase orientation field. Here, we demonstrate that the measurement of colony crystallographic fibers via PLM likewise allow for the back-calculation of the parent grain orientations to two possibilities, from which 24 deterministic colony orientations are forward-calculated. Of these available orientations, we show that four share a axis orientation necessary to satisfy that which PLM is able to measure. We demonstrate this reconstruction method on a synthetic sample where the parent phase microstructure is fully and unambiguously known. Similar to our previous study quantifying the consequences of PLM’s orientation ambiguity in CPFEM predictions (van Wees et al., 2023), we here demonstrate that the choice of one of the four possible orientations leads to significant improvement in deformation field predictions compared to what can be realized via traditional PLM.
@@ 2 Background
In pure titanium and many titanium alloys, the material undergoes an allotropic phase transformation during cooling, transforming from an exclusive body centered cubic structure (the phase), to primarily a hexagonal close packed structure (the phase), with the possibility of some fraction of retained phase (on the order of 10% volume). In pure titanium, the temperature at which this transformation takes place, or the transus temperature, is approximately (Lütjering and Williams, 2003). Generally, as the phase grains (herein: “parent grains”) cool, they tend to transform into multiple distinct sub-grain regions of phase (herein: “ colonies”), which present as individual grains in the cooled microstructure, though collectively retain the geometric morphology of their common parent grain.
@@
Adherent to this relationship, we can calculate the possible colony orientations which may transform from a single parent grain orientation. Described in detail in (Glavicic et al., 2003), the relationship takes the mathematical form:
@@
In (Glavicic et al., 2003), a method is described in which the coupling between the and orientations are used to determine the orientation field of the parent grain structure that likely existed above the transus temperature given a field of measured orientations. This is achieved via the consideration of spatial location/contiguity of measured colonies who have collective intersect among their possible parent variants. Ideally, a collection of contiguous colonies intersect at only one parent variant, allowing for unambiguous determination of the parent grain orientation. This method has been demonstrated on two-dimensional orientation fields (e.g., via electron backscatter diffraction (EBSD) (Glavicic et al., 2003)), as well as three-dimension orientation fields (e.g., via HEDM (Wielewski et al., 2015; Kasemer et al., 2017a)), leading to plausible results regarding the possible orientation field that existed above the transus temperature.
@@ 2.2 Measurement of Fibers via PLM
More recently, it has been shown that the measurement of the field of refracted intensities while rotating a sample comprised of HCP crystals allows for the determination of the field of axis orientations (described in detail in (Jin and De Graef, 2020)). While promising as a method for orientation determination, the prime limitation of PLM in this regard is its inability to determine the full orientation of the crystal. Specifically, PLM is unable to discern:
@@ 2.2.1 Ambiguity of Rotation About the Axis
Discussed in detail in (Jin and De Graef, 2020), the anisotropic nature of optical refraction in HCP materials renders PLM unable to discern the amount of rotation about the axis. This may be best understood via consideration of the Euler-Bunge orientation parameterization (Bunge, 1982): , or rotations about the , , axes, respectively. In the Euler-Bunge parameterization, and determine the orientation of the axis, while determines the rotation about the axis. PLM, thus, is able determine only and , while is left ambiguous.
@@
We can compactly represent the fibers that PLM is able to measure on pole figures. As demonstration, we present a pole figure in Figure 1, depicting the orientations of the axes for the colonies that may descend from a single parent grain with a random orientation (as calculated following the description in Section 2.1). We note the presence of 6 unique peaks (along with 6 symmetrically equivalent negative peaks that appear on the reverse of the pole figure, not shown). This overall indicates that of the 12 possible colony orientations from a single parent grain orientation, two each share the same axis orientation—i.e., two colony orientations lie along the same crystallographic fiber (we provide a demonstration of this property via consideration of the Rodrigues parameterization in A.1).
@@ 2.2.2 Ambiguity of Axis Orientation Due to Periodicity in Measured Intensity
For PLM characterization of HCP specimens, the light intensity data that is collected as the specimen is rotated in relation to the polarization filter displays a periodicity of (as demonstrated in (Jin and De Graef, 2020)). As a result of this periodicity in measured intensities, there exists two axis orientations which cannot be differentiated via PLM 111We note that similar ambiguity exists in ultrasonic characterization methods, particularly spatially resolved acoustic spectroscopy (SRAS), where the observed periodicity is due to the measured directional wave velocity’s proportionality to the anisotropic elastic moduli (Dryburgh et al., 2020).. Inspecting the consequences on the Euler-Bunge angles, assuming that the sample in question is rotated about the sample direction during data collection, the periodicity affects the measurement of —the measured angle will either be correct, or potentially a mis-indexing of the true orientation of .
@@ 3 Methodology for Reconstruction of Quasi-Determinstic Orientations
Having established that PLM is able to measure colony fibers as opposed to unambiguous orientations, we now present a methodology to reconstruct the colony orientations to one of four possibilities—i.e., a quasi-deterministic approach. Our methodology is inspired by the approach utilized in (Glavicic et al., 2003) in which the parent grain orientations likely present above the transus temperature may be calculated utilizing the deterministic colony orientations. Overall, the flow of our method is summarized as follows:
@@ 3.1.2 Clustering Colonies Based on Shared Lineage
Regarding the former, we assume that colonies that share a common parent grain lineage are spatially contiguous. To efficiently test for spatial contiguity of colonies, we construct a graph representation of a polycrystal, where the colonies in a polycrystal are represented as “nodes” in the graph, and the graph “edges” describe the connection between nodes, ultimately providing a simplified spatial map of the polycrystal (Pagan et al., 2022). Identification of colonies and their neighbors to facilitate construction of the graph follows standard grain indexing practices, such as with EBSD data. We represent the graph of the polycrystal as a (generally sparse) adjacency matrix, whose row and column indices refer to the indexed IDs of the colonies, and entries of “1” indicate connectivity or sharing a graph edge (i.e., a neighboring / spatially contiguous colony), and entries of “0” indicate a non-neighboring colony, allowing for efficient identification of neighborhood.
@@
Algorithmically, we start at any point of the graph and work neighbor-to-neighbor through the graph in a routine “flood fill” scheme (Wielewski et al., 2015). In essence, we work through the graph testing (naturally) spatial contiguity of the colonies, as well as whether each neighboring colony shares a common parent grain lineage based on the above-described criteria comparing the angle between the axes. If two colonies satisfy both of these criteria, we record that they must share the same parent grain ID. After completely filling through the graph, we are left with a field of parent grain IDs assigned to each colony, or (in other words) networks of colonies belonging to the same parent grain. Further, for each parent grain, we generate a pole figure considering all of the measured and reflected axes for every colony determined to share the same parent grain lineage. We normalize the pole figure such that the integrated surface intensity is 1. We use this pole figure to register against a dictionary of pole figures to determine the parent grain orientations.
@@ 3.2 Registration of Measured Pole Figure Against Dictionary Entries
We next generate a set of idealized pole figures representing the possible true and reflected axes that descend from parent grains of known orientation, termed a ‘dictionary’. To achieve this, we start by generating a list of parent grain orientations. We utilize the nodes of a finite element mesh of the cubic symmetry fundamental region of Rodrigues space (Kumar and Dawson, 1998) to generate a list of unique orientations (i.e., no symmetrically equivalent orientations) with a dense packing across the fundamental region (i.e., providing fine resolution of the parent grain orientations). For each of these parent grain orientations, we calculate the orientation of the axes of their colony variants as well as their reflections. We then construct a pole figure based on the set of unique axes, assuming a surface Gaussian distribution of minimal spread (average ) at each point to later aid in registration (this allows for some mismatch in registration in case of potential misaligned axes, due to imperfect transformation or uncertainty in measurement). Again, we normalize the pole figures such that the integrated surface intensity is 1. Figure 2 depicts a typical dictionary entry (itself a summation of the true and reflected pole figures depicted previously in Figure 1).
@@ 3.3 Forward Calculation of Quasi-Determinstic Colony Orientations
Knowing the two parent grain orientations which result in the same orientations of colony axes and their reflections, we calculate the true colony orientations based on the Burgers orientation relationship. For each of the two parent grain orientations, we calculate 12 unique colony orientations (24 total). We can show that among the 12 orientations from each grain, two each share the same axis orientation and cannot be further distinguished via PLM (see: Appendix A.1). Further, there exist two reflected orientations that share the reflected axis orientation of the measured, which again we cannot distinguish between via PLM. We are thus able to narrow down the possible colony orientations to one of four values.

### v2

@@ 1 Introduction
Research into fundamental crystal-scale behavior has led to a decrease in uncertainty with respect to macroscopic mechanical behavior of polycrystalline materials, which has in turn facilitated the design of materials and engineering components which are able to meet increasingly stringent demands in modern machinery. Significant experimental developments were made in the latter half of the 20th century regarding a deeper understanding of how the morphology of grains and the collective distribution of crystallographic orientations (or: crystallographic texture) influence the behavior of materials, primarily in terms of the yield strength and ductility. In parallel, developments in microstructurally-sensitive deformation modeling techniques have similarly enjoyed significant progress. Various techniques have been developed in this regard, from mean field methods in the 1930s [1, 2], to self-consistent modeling primarily developed in the 1980s and 1990s [3]. First developed in the late 1980s [4, 5] as a method for modeling texture evolution, crystal plasticity finite element modeling (CPFEM) has evolved to a full-field method capable of capturing the intragrain deformation response of high-fidelity representations of microstructures (considering both crystallographic texture and explicit representations of complex grain morphology) [6, 7, 8], and has been proven adept at the prediction of both intragrain and macroscopic behavior, and has offered unparalleled insight into the relationship of elasticity and plasticity at those scales (and thus our best window into structure-property relationships).
@@
In all of the above modeling methods, the models require some degree of experimental material characterization or micromechanical testing to aid in initiation, verification, or validation of the model. For CPFEM specifically, proper instantiation of a virtual polycrystalline sample that faithfully represents a physical sample requires (at least) information about the geometric morpohology of the grains, as well as the grains orientations (i.e., crystallographic texture) [8]. Most often, two-dimensional characterization methods such as electron backscatter diffraction (EBSD) are used to measure both granular morphology and orientations on the surfaces of samples, and various methods have been developed to infer the three-dimensional morphology (either statistically [9, 10], based on oblique surface measurements [11], or destructively in the case of serial sectioning [12]). More recently, high energy diffraction microscopy (HEDM) has offered the ability to non-destructively map orientations in three dimensions [13, 14, 15, 16, 17], which can be used to directly construct digital twins of actual physical samples [18], rather than statistically-representative virtual samples. Further, in the case of HEDM, micromechanical testing may be performed on the samples after initial characterization, allowing for unprecedented experimental deduction of structure-property relationships [19, 20].
@@
Unfortunately, while EBSD and HEDM have become well-accepted as characterization techniques and are indeed powerful tools which have progressed significantly in terms of cost and ease of use, they suffer from practical limitations. Both are relatively time consuming (e.g., PLM can image large scans at a sub micron resolution in a fraction of the time—minutes versus hours—compared to EBSD or HEDM), expensive (e.g., EBSD is generally charged by the hour, and is thus naturally more expensive, while HEDM requires a synchrotron light source), and require a relatively high degree of training necessary to produce skilled researchers capable of performing proper, meaningful measurements compared to the method of PLM (this latter point is not meant to disparage these techniques, but rather highlight the inherent difficulty in the theory necessary to expertly employ them). Consequently, it is beneficial to develop methods which significantly reduce cost, lead-time, and/or the knowledge/skill barrier necessary to produce adequate characterization results, especially in cases where high-throughput is desired. Recently, significant research attention has been given to the development of ultrasonic [21, 22, 23, 24] and optical techniques in an effort to address these problems. Of interest to this study, we highlight the advent of polarized light microscopy (PLM) as a characterization method for materials comprised of crystals exhibiting hexagonal symmetry [25, 26, 27]. PLM offers the ability to measure spatial fields of (partial) crystal orientations both relatively rapidly and using relatively affordable equipment (compared to EBSD), as well as with fewer material preparation steps. However, PLM is limited in that it only measures an incomplete description of the orientation of HCP crystals, which limits its applicability as a characterization method for micromechanical simulations owing to the potential for high errors in predictions [28].
@@
In this study, we outline a method to reconstruct full orientations of the crystals in titanium materials to one of four orientations—i.e., a quasi-deterministic approach. As prime inspiration for this study, we consult the work of Glavicic et. al. [29], which established a method exploiting the physics of the allotropic phase transformation that many titanium alloys undergo during cooling in an effort to deduce the orientation field of the BCC ( phase) microstructure that likely existed at high temperature based on the measurement of the room temperature phase orientation field. Here, we demonstrate that the measurement of colony crystallographic fibers via PLM likewise allow for the back-calculation of the parent grain orientations to two possibilities, from which 24 deterministic colony orientations are forward-calculated. Of these available orientations, we show that four share a axis orientation necessary to satisfy that which PLM is able to measure. We demonstrate this reconstruction method on a synthetic sample where the parent phase microstructure is fully and unambiguously known (we opt for virtual samples to unambiguously validate our method rather than comparison to another method, which is prone to experimental uncertainty as well as sensitivity to reconstruction parameters). Similar to our previous study quantifying the consequences of PLM’s orientation ambiguity in CPFEM predictions [28], we here demonstrate that the choice of one of the four possible orientations leads to significant improvement in deformation field predictions compared to what can be realized via traditional PLM.
@@ 2 Background
In pure titanium and many titanium alloys, the material undergoes an allotropic phase transformation during cooling, transforming from an exclusive body centered cubic structure (the phase), to primarily a hexagonal close packed structure (the phase), with the possibility of some fraction of retained phase (on the order of 10% volume). In pure titanium, the temperature at which this transformation takes place, or the transus temperature, is approximately [30]. Generally, as the phase grains (herein: “parent grains”) cool, they tend to transform into multiple distinct sub-grain regions of phase (herein: “ colonies”), which present as individual grains in the cooled microstructure, though collectively retain the approximate geometric morphology of their common parent grain (or, at least, the collection of colonies which descend from a common parent grain will be connected in a spatially contiguous network, if not perfectly adhering to the exact morphology of the parent grain).
@@
Adherent to this relationship, we can calculate the possible colony orientations which may transform from a single parent grain orientation. Described in detail in [29], the relationship takes the mathematical form:
@@
In [29], a method is described in which the coupling between the and orientations are used to determine the orientation field of the parent grain structure that likely existed above the transus temperature given a field of measured orientations. This is achieved via the consideration of spatial location/contiguity of measured colonies who have collective intersect among their possible parent variants. Ideally, a collection of contiguous colonies intersect at only one parent variant, allowing for unambiguous determination of the parent grain orientation. This method has been demonstrated on two-dimensional orientation fields (e.g., via electron backscatter diffraction (EBSD) [29]), as well as three-dimension orientation fields (e.g., via HEDM [31, 32]), leading to plausible results regarding the possible orientation field that existed above the transus temperature.
@@ 2.2 Measurement of Fibers via PLM
More recently, it has been shown that the measurement of the field of refracted intensities while rotating a sample comprised of HCP crystals allows for the determination of the field of axis orientations (described in detail in [25]). While promising as a method for orientation determination, the prime limitation of PLM in this regard is its inability to determine the full orientation of the crystal. Specifically, PLM is unable to discern:
@@ 2.2.1 Ambiguity of Rotation About the Axis
Discussed in detail in [25], the anisotropic nature of optical refraction in HCP materials renders PLM unable to discern the amount of rotation about the axis. This may be best understood via consideration of the Euler-Bunge orientation parameterization [33]: , or rotations about the , , axes, respectively. In the Euler-Bunge parameterization, and determine the orientation of the axis, while determines the rotation about the axis. PLM, thus, is able determine only and , while is left ambiguous.
@@
We can compactly represent the fibers that PLM is able to measure on pole figures. As demonstration, we present a pole figure in Figure 1, depicting the orientations of the axes for the colonies that may descend from a single parent grain with a random orientation (as calculated following the description in Section 2.1). We note the presence of 6 unique peaks (along with 6 symmetrically equivalent negative peaks that appear on the reverse of the pole figure, not shown). This overall indicates that of the 12 possible colony orientations from a single parent grain orientation, two each share the same axis orientation—i.e., two colony orientations lie along the same crystallographic fiber (we provide a demonstration of this property via consideration of the Rodrigues parameterization in Appendix A.1).
@@ 2.2.2 Ambiguity of Axis Orientation Due to Periodicity in Measured Intensity
For PLM characterization of HCP specimens, the light intensity data that is collected as the specimen is rotated in relation to the polarization filter displays a periodicity of (as demonstrated in [25]). As a result of this periodicity in measured intensities, there exists two axis orientations which cannot be differentiated via PLM 111We note that similar ambiguity exists in ultrasonic characterization methods, particularly spatially resolved acoustic spectroscopy (SRAS), where the observed periodicity is due to the measured directional wave velocity’s proportionality to the anisotropic elastic moduli [22].. Inspecting the consequences on the Euler-Bunge angles, assuming that the sample in question is rotated about the sample direction during data collection, the periodicity affects the measurement of —the measured angle will either be correct, or potentially a mis-indexing of the true orientation of .
@@ 3 Methodology for Reconstruction of Quasi-Determinstic Orientations
Having established that PLM is able to measure colony fibers as opposed to unambiguous orientations, we now present a methodology to reconstruct the colony orientations to one of four possibilities—i.e., a quasi-deterministic approach. Our methodology is inspired by the approach utilized in [29] in which the parent grain orientations likely present above the transus temperature may be calculated utilizing the deterministic colony orientations. Overall, the flow of our method is summarized as follows:
@@ 3.1.2 Clustering Colonies Based on Shared Lineage
Regarding the former, we assume that colonies that share a common parent grain lineage are spatially contiguous. To efficiently test for spatial contiguity of colonies, we construct a graph representation of a polycrystal, where the colonies in a polycrystal are represented as “nodes” in the graph, and the graph “edges” describe the connection between nodes, ultimately providing a simplified spatial map of the polycrystal [34]. Identification of colonies and their neighbors to facilitate construction of the graph follows standard grain indexing practices, such as with EBSD data. We represent the graph of the polycrystal as a (generally sparse) adjacency matrix, whose row and column indices refer to the indexed IDs of the colonies, and entries of “1” indicate connectivity or sharing a graph edge (i.e., a neighboring / spatially contiguous colony), and entries of “0” indicate a non-neighboring colony, allowing for efficient identification of neighborhood.
@@
Algorithmically, we start at any point of the graph and work neighbor-to-neighbor through the graph in a routine “flood fill” scheme [31]. In essence, we work through the graph testing (naturally) spatial contiguity of the colonies, as well as whether each neighboring colony shares a common parent grain lineage based on the above-described criteria comparing the angle between the axes. If two colonies satisfy both of these criteria, we record that they must share the same parent grain ID. After completely filling through the graph, we are left with a field of parent grain IDs assigned to each colony, or (in other words) networks of colonies belonging to the same parent grain. Further, for each parent grain, we generate a pole figure considering all of the measured and reflected axes for every colony determined to share the same parent grain lineage. We normalize the pole figure such that the integrated surface intensity is 1. We use this pole figure to register against a dictionary of pole figures to determine the parent grain orientations.
@@ 3.2 Registration of Measured Pole Figure Against Dictionary Entries
We next generate a set of idealized pole figures representing the possible true and reflected axes that descend from parent grains of known orientation, termed a ‘dictionary’. To achieve this, we start by generating a list of parent grain orientations. We utilize the nodes of a finite element mesh of the cubic symmetry fundamental region of Rodrigues space [35] to generate a list of unique orientations (i.e., no symmetrically equivalent orientations) with a dense packing across the fundamental region (i.e., providing fine resolution of the parent grain orientations). For each of these parent grain orientations, we calculate the orientation of the axes of their colony variants as well as their reflections. We then construct a pole figure based on the set of unique axes, assuming a surface Gaussian distribution of minimal spread (average ) at each point to later aid in registration (this allows for some mismatch in registration in case of potential misaligned axes, due to imperfect transformation or uncertainty in measurement). Again, we normalize the pole figures such that the integrated surface intensity is 1. Figure 2 depicts a typical dictionary entry (itself a summation of the true and reflected pole figures depicted previously in Figure 1).
@@ 3.3 Forward Calculation of Quasi-Determinstic Colony Orientations
Knowing the two parent grain orientations which result in the same orientations of colony axes and their reflections, we calculate the true colony orientations based on the Burgers orientation relationship. For each of the two parent grain orientations, we calculate 12 unique colony orientations (24 total). We can show that among the 12 orientations from each grain, two each share the same axis orientation and cannot be further distinguished via PLM (see: A.1). Further, there exist two reflected orientations that share the reflected axis orientation of the measured, which again we cannot distinguish between via PLM. We are thus able to narrow down the possible colony orientations to one of four values.
3.4 Experimental Complications
We focus this study on the theoretical establishment of the technique presented above, and (to be discussed in Section 5) the consequences when utilized in crystal plasticity finite element simulations, rather than on the experimental method of polarized light microscopy. There exists, however, some practical hindrances in terms of the experimental employ of this method that must be addressed. First, we note that an experimental sample must have proper surface preparation—similar to EBSD—such that the signal to noise ratio is maximized, providing confidence in the measurement of the field of axes. Failure to properly prepare the surface of the sample will increase uncertainty in (partial) orientation indexing, thus increasing uncertainty in the method presented in this study. Second, we note that we have thus far assumed that each colony variant for a given parent grain is present, such that the determination of the parent grains is limited to two orientations. If only one colony orientation from a given parent grain is present, then the parent grain orientation can only be determined to a fiber. This quickly collapses to only two orientations—particularly when considering the necessary reflections of the colonies. While we do not envision that there will exist a large fraction of parent grains which transform wholly to a single colony, there is a larger probability of edge cases—particularly along the physical boundary of the measured sample—where only a single colony variant from a parent grain is present, and thus confidence along the boundary of the sample measurement will be lower. Finally, we acknowledge that the Burgers orientation relationship may lead to imperfect transformations (i.e., of Equation 2 may vary slightly), which may further introduce uncertainty into the reconstruction.
Collectively, an experimental implementation would likely be accompanied with a confidence metric to quantify the certainty of the orientation reconstruction. Motivated primarily by the facts that measurements may contain noise, and that transformations may be imperfect, we note that the pole figures constructed from measured poles and their reflections (similar to that presented in Figure 2) may contain peaks that are not perfect “points” on the pole figure, but rather “spreads”. Since the dictionary entries are likewise constructed from peaks with a nominal degree of spread, this allows for imperfect mismatch between what is measured and what exists in the dictionary. We envision that a confidence metric could be constructed locally (i.e., as a field), and quantify (for example) the degree of mismatch between the measured peaks at that location and those in the chosen dictionary entry, likely via a misorientation metric. Other similar metrics could be borrowed from EBSD and appropriately modified for the case here. We do not pursue the construction of such a metric in this study, as we focus instead on the establishment and verification of the method via the idealized theoretical problem (as will be further motivated in Section 4).